Half of Haley's age is 6 years less than her age 9 years ago. How old is Haley?
Answer by stanbon(75887)

Half of Haley's age is 6 years less than her age 9 years ago.
How old is Haley?
-----Equation:
(1/2)h = (h-9)-6
----
(1/2)h = h-15
h = 2h-30
h = 30 yrs (Haley's age now)
=============================
Cheers,
Stan H.
=========